OpenStudy (anonymous):

Use the method of Lagrange multipliers to optimize the function subject to the given constraint. f(x, y) = 4x + 6y - 2x2 - 7y2 where 3x + 9y = 9 Find the coordinates of the optimal point subject to the constraint.

OpenStudy (amistre64):

ok , where are you getting stuck at with this?

OpenStudy (anonymous):

i need the value for x and the value for y and f(x,y) but I dont know where to start...

OpenStudy (amistre64):

well, we can start by creating a single function from these F(x,y,L). Either by adding or subtracting the consttraint

OpenStudy (amistre64):

like this: F(x, y,L) = 4x + 6y - 2x2 - 7y2 + L(3x + 9y - 9)

OpenStudy (amistre64):

that is the start :) we can distribute the L thru the () .... i used L to represent the Lagrange multiplier

OpenStudy (anonymous):

okk thanks is L lambda?

OpenStudy (amistre64):

yes, L = lambda :) F(x,y,L) = 4x + 6y - 2x2 - 7y2 + L3x + L9y - L9 Now we find our partials: Fx Fy FL

OpenStudy (anonymous):

how would you use the close point method to classify the optimal point. to find out if the optimal point is a max or min?

OpenStudy (amistre64):

close point sounds like you find point near it and determine whether they rise of fall

OpenStudy (amistre64):

i dont think ive ever actually tried the "method" but thats what it sounds like

OpenStudy (anonymous):

oh ok sounds good thanks for your help ! im gona try to see if i can figure it out now :/

OpenStudy (amistre64):

Fx = 4x -2x2 + L3x = 4 -4x + L3 = 0; solve for x Fy = 6y -7y2 +L9y = 6 - 14y +L9 = 0; solve for y FL = L3x+L9y-L9 = 3x +9y -9 = 0; substitute in x and y and solve for L

OpenStudy (anonymous):

ok thank you so so much!

OpenStudy (amistre64):

let me know what you get for an answer if you want :) and i can dbl chk it